OpenStudy (anonymous):

4x^2+9y^2-8x-18y-23=0 find the center of this ellipse

OpenStudy (anonymous):

\[\frac{ (x-h)^2 }{ a^2 }+\frac{( y-k)^2 }{ b^2}\]

OpenStudy (anonymous):

Thats the formula

Nnesha (nnesha):

\[\frac{ (x-h)^2 }{ a^2 }+\frac{( y-k)^2 }{ b^2}=1\] now looks good

OpenStudy (anonymous):

Yea

OpenStudy (anonymous):

enjoy completing the square twice raft of algebra

OpenStudy (anonymous):

:(

OpenStudy (anonymous):

I got 1,1

Nnesha (nnesha):

 x's and y's now apply completing the square method for both parenthses